第38回環論および表現論シンポジウム（2005年）プログラム

９月２日（金曜日) ９：４５−１０：３０ 大貫洋介（東京農工大学） Stable equivalences related with syzygy functors １０：４５−１１：３０ 長瀬潤（奈良工業高等専門学校） Hochschild cohomology of stratified algebras １１：４５−１２：３０ 西田憲司（信州大学） Linkage and duality of modules １３：４５−１４：４５ 有木進（京都大学） Algebras arising from Schur-Weyl type dualities (I) １５：００−１５：４５ 柳川浩二（大阪大学） Castelnuovo-Mumford regularity for complexes and weakly Koszul modules １６：００−１６：４５ 本瀬香（弘前大学） Integral group algebras and cyclotomic polynomials １７：００−１８：００ Raphaël Rouquier （Université Denis Diderot - Paris 7） Broué’s conjecture: methods and results (I)

９月３日 (土曜日) ９：４５−１０：３０ 伊山修（兵庫県立大学） Orthogonality of subcategories １０：４５−１１：３０ 高橋亮（明治大学) A generalization of n-torsionfree modules １１：４５−１２：３０ 越谷重夫（千葉大学） Broué’s abelian defect group conjecture

１３：４５−１４：４５ 有木進（京都大学) Algebras arising from Schur-Weyl type dualities (II)

１５：００−１６：００ Raphaël Rouquier （Université Denis Diderot - Paris 7） Broué’s conjecture: methods and results (II)

１６：１５−１７：００ 太刀川弘幸（筑波大学） QF rings and QF associated graded rings

９月４日（日曜日）

９：４５−１０：４５ Raphaël Rouquier (Université Denis Diderot - Paris 7) Broué’s conjecture: methods and results (III)

１１：００−１２：００ 有木進（京都大学) Algebras arising from Schur-Weyl type dualities (III)